OpenStudy (anonymous):

Use the commutative property to rewrite the expression: negative 8 x times 7 y

OpenStudy (anonymous):

lots of choices here i one \[-8x\times 7y=-8xy\times 7\] or \[-8\times 7xy\] is another

OpenStudy (anonymous):

i.e. change the order of the symbols any way you lik e